OpenStudy (anonymous):

A bag contains 7 green marbles, 9 red marbles, 10 orange marbles, 5 brown marbles, and 10 blue marbles. You choose a marble, replace it, and choose again. Find P(red, then blue).

OpenStudy (anonymous):

To find the answer of this problem, you have to multiply the probability of choosing red by the probability of choosing blue. P(red)= 9/41 (blue)=10/41 9/41 (10/41) = 0.0535 The probability is about 5.35%